If we choose velocity V, acceleration A and forceF as fundamental quantities then the angular momentum in terms of V,A and F would be ??
ML2T-1
FA-1V
FV3A-2
FV2A-1

Grade:12

2 Answers

Arun
4 years ago
let the dimensional formula of mass is M = EaFbVc
 E=ML2T-2  ,  F =MLT-2  , V=LT-1
on putting these
    M = Ma+b L2a+b+c T-2a-2b-c
equating powers ,we get
 a=1 , b=0, c=-2
 dimensional formula of mass= E1.F0 V-2
 
Khimraj
4 years ago

D) ML² T-1

 

 

Explanation:

angular momentum = moment of inertia * angular velocity

angular momentum = mass *radius² *time inverse

which means its fundamental value will be ML²T-1